3. (25 points) Let TA and Te be two function returning the running time of algorithms A and B, defined by the recusions TA(n) T(2 and TR(n) T(2n2. Find the largest integer value of a for which algorithm B is asymptotically faster than A. Show your work and justify your answer in the PDF file. Answer in the MyCourses quiz.
Answer #1

TA=7TA(n/2)+n2.

Applying master's methon , solution to this recurrence is : O(nlog27).

Now, consider the second recurrence:

T_B=\alpha T_B(n/4)+n^2.

Now for value of \alpha<16 , solution to this recurrence will be O(n2).

for value of \alpha=16 , solution to this recurrence will be O(n2log2n).

for value of \alpha>16 , solution to this recurrence will be O(n^{log_4\alpha}) , which is asymptotically not faster than

O(nlog27).

So, 16 is the largest integer for \alpha , so that B is faster than A.